The sum of a number and 4 times its reciprocal is 25 over 6...

equation is: X+4(1/X)+25/6... but how do I solve this?
Thank you!

Answer by checkley71(8403)

X+4/X=25/6 OR (X~2+4)/X=25/6 OR 6X~2+24=25X OR 6X~2-25X=24=0 OR
(3X-8)(2X-3)=0 OR 3X-8=0 OR 3X=8 OR X=8/3 AND 2X-3=0 OR 2X=3 OR X=2/3
PROOF 8/3+4/(8/3)=25/6 OR 8/3+4*(3/8)=25/6 OR 8/3+12/8=25/6 OR
64/24+36/24=25/6 OR 100/24=25/6 OR 25/6=25/6